# Score–Curvature Out-of-distribution Proximity Evaluator for Diffusion

### v1

@@ 2.1 Score-Based Diffusion Models
Diffusion models (Sohl-Dickstein et al., 2015; Song et al., 2021) train a neural network to approximate the score function $s_{\theta}(x_{t},t)\approx\nabla_{x}\log p_{t}(x_{t})$ of progressively noised data $x_{t}$. At step $t$, Gaussian noise with variance $\sigma_{t}^{2}$ is added according to a schedule that grows with $t$, defining a distribution over noise scales from which training samples are drawn so the network learns to denoise across the full noise range. Given a noised input, the network predicts either the score or an equivalent representation (e.g. denoised data or noise), from which the score can be recovered; see Appendix B for details. Sampling then integrates the corresponding reverse dynamics to map Gaussian noise back into data. Early noise levels preserve fine detail, while later ones retain only coarse structure.
DDPM and EDM are discrete and continuous instances of this shared framework: both ultimately train a network to approximate $\nabla_{x}\log p_{t}(x_{t})$. This common structure underlies our OOD detection method. For completeness, we provide a more detailed derivation of DDPM, denoising score matching, and EDM in Appendix B.
@@ 2.3 The Score-Curvature Ratio
A final subtlety is that the typicality ratio admits a global sign ambiguity: the squared norm in the numerator discards orientation, while the Jacobian trace in the denominator retains sign information. We correct this with a global sign factor, as detailed in Appendix C, and confirm in Appendix I that this correction is necessary for stable OOD performance in vision-based tasks.
@@ 3.3 Offline Selection of Noise Levels for SCOPED
In contrast, for our vision experiments, we estimate the signal-to-noise ratio (SNR) of the distribution $p_{t}(x_{t})$ offline using in-distribution data (cf. Appendix D), at different points along the forward diffusion process $t$. SNR decays monotonically: early timesteps preserve fine detail, while later ones approach noise dominance. To ensure robustness across this spectrum, we evaluate SCOPED at two points: an early step ($t=1$) that maximizes fine information, and a mid-level step ($t=300$) where about 95% of the signal remains, the SNR curve enters a roughly linear decline, and coarse structure is retained while some fine detail is suppressed.
@@ 5.1 Reinforcement Learning Case Studies
On DeepMind Control Suite tasks, SCOPED achieves perfect separation across all tested shifts (Appendix F, Table 2). These results validate that the method captures distributional divergence not only in cases where the state–action–reward distributions are clearly non-overlapping, such as humanoid-stand versus humanoid-walk (cf. Figure 1), but also in settings where overlap is expected, such as finger-turn or reacher tasks. In these latter cases, the same dynamics and similar reward structures generate transitions that occupy overlapping regions of the state-action space, yet SCOPED still produces strong separation and perfect out-of-distribution detection accuracy. This outcome would not have been obvious a priori and suggests that the statistic is sensitive to more subtle distributional cues than reward specification alone. While these experiments may appear straightforward a posteriori, they serve as valuable sanity checks: they confirm that SCOPED behaves in line with intuition when distributions are clearly distinct, while also demonstrating sensitivity in cases where overlap exists and separability is less obvious. Together, they provide a foundation before turning to more challenging benchmarks.
@@ Results.
With this setup, the results show that SCOPED achieves competitive or state-of-the-art detection accuracy while requiring orders of magnitude fewer effective function evaluations than prior methods. The default two-step variant (2F+2J) offers robustness across dataset pairs while remaining far cheaper than trajectory-based approaches. Ablations over different early–mid step pairs (Appendix H) confirm that AUROC remains consistently high across choices, demonstrating that the two-step procedure provides robust performance across timesteps.
We also evaluate SCOPED (Single), which fixes the noise level at $t=300$ chosen once from in-distribution data. By omitting the earlier fine-detail step, this variant sacrifices some robustness, yet it demonstrates that strong performance is achievable with only 1F+1J. Ablations varying the mid-step choice (Appendix G) confirm that performance remains relatively insensitive to the exact choice of step.
@@ 6 Conclusion
Reproducibility Statement
We have taken several steps to ensure the reproducibility of our results. Implementation details for SCOPED are provided in Appendix E, with full descriptions of noise level selection in Section 3.3, score computation in Appendix B, and sign correction in Appendix C. Additional ablations, including single-step versus two-step variants and noise-level sensitivity, are reported in Appendix G and Appendix H. Vision experiments use the open-source CelebA DDPM checkpoint from Heng et al. (2024), RL experiments rely on the standard D4RL datasets and our own trained diffusion models. We will release code for the vision experiments to ensure easy replication.
@@ Appendix A Usage of Large Language Models (LLMs)
Appendix B Additional Background on Diffusion Models
@@
Appendix C Sign Ambiguity in the Typicality Ratio
@@
Our implementation applies this factor to all OOD scores. We show in Appendix I that this correction is important for vision tasks.
Appendix D Signal-to-Noise Ratio of CIFAR-10
@@
Appendix E Implementation Walkthrough (Vision Experiments)
@@
    Select noise levels. For vision tasks, we evaluate at two timesteps: $t=1$ (fine detail preserved) and $t=300$ (mid-level, $\sim$95% signal retained), chosen from the signal-to-noise ratio curve (Appendix D).
@@
    and apply a global sign factor as described in Appendix C, where $\varepsilon>0$ is a small constant added for numerical stability to avoid division by zero.
@@
Appendix F Offline DMC Results
@@
Appendix G Ablation Over Single Step Choices in Noise Schedule for SCOPED
@@
Appendix H Ablation Over Two Step Choices in Noise Schedule for SCOPED
@@
Appendix I Ablation Over Two Step Choices in Noise Schedule for SCOPED without Sign Ambiguity Term
We repeat the two-step ablation without applying the global sign correction described in Appendix C. Table 5 shows that while AUROC remains non-trivial when SCOPED is applied without a sign correction, average performance drops to 0.81 compared to 0.84. This confirms that the sign term is necessary for stable performance in vision benchmarks, even though raw score–curvature ratios can still capture some separation.

### v2

@@ 2.1 Score-Based Diffusion Models
Diffusion models (Sohl-Dickstein et al., 2015; Song et al., 2021) train a neural network to approximate the score function $s_{\theta}(x_{t},t)\approx\nabla_{x}\log p_{t}(x_{t})$ of progressively noised data $x_{t}$. At step $t$, Gaussian noise with variance $\sigma_{t}^{2}$ is added according to a schedule that grows with $t$, defining a distribution over noise scales from which training samples are drawn so the network learns to denoise across the full noise range. Given a noised input, the network predicts either the score or an equivalent representation (e.g. denoised data or noise), from which the score can be recovered; see Appendix C for details. Sampling then integrates the corresponding reverse dynamics to map Gaussian noise back into data. Early noise levels preserve fine detail, while later ones retain only coarse structure.
DDPM and EDM are discrete and continuous instances of this shared framework: both ultimately train a network to approximate $\nabla_{x}\log p_{t}(x_{t})$. This common structure underlies our OOD detection method. For completeness, we provide a more detailed derivation of DDPM, denoising score matching, and EDM in Appendix C.
@@ 2.3 The Score-Curvature Ratio
A final subtlety is that the typicality ratio admits a global sign ambiguity: the squared norm in the numerator discards orientation, while the Jacobian trace in the denominator retains sign information. We correct this with a global sign factor, as detailed in Appendix D, and confirm in Appendix J that this correction is necessary for stable OOD performance in vision-based tasks.
@@ 3.3 Offline Selection of Noise Levels for SCOPED
In contrast, for our vision experiments, we estimate the signal-to-noise ratio (SNR) of the distribution $p_{t}(x_{t})$ offline using in-distribution data (cf. Appendix E), at different points along the forward diffusion process $t$. SNR decays monotonically: early timesteps preserve fine detail, while later ones approach noise dominance. To ensure robustness across this spectrum, we evaluate SCOPED at two points: an early step ($t=1$) that maximizes fine information, and a mid-level step ($t=300$) where about 95% of the signal remains, the SNR curve enters a roughly linear decline, and coarse structure is retained while some fine detail is suppressed.
@@ 5.1 Reinforcement Learning Case Studies
On DeepMind Control Suite tasks, SCOPED achieves perfect separation across all tested shifts (Appendix G, Table 2). These results validate that the method captures distributional divergence not only in cases where the state–action–reward distributions are clearly non-overlapping, such as humanoid-stand versus humanoid-walk (cf. Figure 1), but also in settings where overlap is expected, such as finger-turn or reacher tasks. In these latter cases, the same dynamics and similar reward structures generate transitions that occupy overlapping regions of the state-action space, yet SCOPED still produces strong separation and perfect out-of-distribution detection accuracy. This outcome would not have been obvious a priori and suggests that the statistic is sensitive to more subtle distributional cues than reward specification alone. While these experiments may appear straightforward a posteriori, they serve as valuable sanity checks: they confirm that SCOPED behaves in line with intuition when distributions are clearly distinct, while also demonstrating sensitivity in cases where overlap exists and separability is less obvious. Together, they provide a foundation before turning to more challenging benchmarks.
@@ Results.
With this setup, the results show that SCOPED achieves competitive or state-of-the-art detection accuracy while requiring orders of magnitude fewer effective function evaluations than prior methods. The default two-step variant (2F+2J) offers robustness across dataset pairs while remaining far cheaper than trajectory-based approaches. Ablations over different early–mid step pairs (Appendix I) confirm that AUROC remains consistently high across choices, demonstrating that the two-step procedure provides robust performance across timesteps.
We also evaluate SCOPED (Single), which fixes the noise level at $t=300$ chosen once from in-distribution data. By omitting the earlier fine-detail step, this variant sacrifices some robustness, yet it demonstrates that strong performance is achievable with only 1F+1J. Ablations varying the mid-step choice (Appendix H) confirm that performance remains relatively insensitive to the exact choice of step.
@@ 6 Conclusion
7 Acknowledgments
This work was supported by the National Science Foundation under Grant No. 2409535.
@@ Appendix A Usage of Large Language Models (LLMs)
Appendix B Reproducibility Statement
We have taken several steps to ensure the reproducibility of our results. Implementation details for SCOPED are provided in Appendix F, with full descriptions of noise level selection in Section 3.3, score computation in Appendix C, and sign correction in Appendix D. Additional ablations, including single-step versus two-step variants and noise-level sensitivity, are reported in Appendix H and Appendix I. Vision experiments use the open-source CelebA DDPM checkpoint from Heng et al. (2024), RL experiments rely on the standard D4RL datasets and our own trained diffusion models. We will release code for the vision experiments to ensure easy replication.
Appendix C Additional Background on Diffusion Models
@@
Appendix D Sign Ambiguity in the Typicality Ratio
@@
Our implementation applies this factor to all OOD scores. We show in Appendix J that this correction is important for vision tasks.
Appendix E Signal-to-Noise Ratio of CIFAR-10
@@
Appendix F Implementation Walkthrough (Vision Experiments)
@@
    Select noise levels. For vision tasks, we evaluate at two timesteps: $t=1$ (fine detail preserved) and $t=300$ (mid-level, $\sim$95% signal retained), chosen from the signal-to-noise ratio curve (Appendix E).
@@
    and apply a global sign factor as described in Appendix D, where $\varepsilon>0$ is a small constant added for numerical stability to avoid division by zero.
@@
Appendix G Offline DMC Results
@@
Appendix H Ablation Over Single Step Choices in Noise Schedule for SCOPED
@@
Appendix I Ablation Over Two Step Choices in Noise Schedule for SCOPED
@@
Appendix J Ablation Over Two Step Choices in Noise Schedule for SCOPED without Sign Ambiguity Term
We repeat the two-step ablation without applying the global sign correction described in Appendix D. Table 5 shows that while AUROC remains non-trivial when SCOPED is applied without a sign correction, average performance drops to 0.81 compared to 0.84. This confirms that the sign term is necessary for stable performance in vision benchmarks, even though raw score–curvature ratios can still capture some separation.